OpenStudy (anonymous):

An urn contains 10 white balls and 5 blue balls. Draws are made repeatedly from the urn as follows. On each draw, a ball is drawn and its color noted; then it is replaced in the urn along with 3 more balls of its color.Find the chance that the first ball drawn is blue, given that the second ball drawn is blue.

OpenStudy (anonymous):

required probability=5/15*8/18 =1/3*4/9=4/27

OpenStudy (anonymous):

Find the chance that the second ball drawn is blue = 5/15*8/18